I need help please: Question is--
Take indicated quadratic inequality x^2 - 13X + 36 < or equal to 0.
A) what type of graph
B) whether it goes down or up
C) Determine size
D) Graph equation
E) Find Vertex
F) Find X intercept
G) Find Y intercept
H) what's the focal point

Answer by richwmiller(17219)

(x-9) (x-4)<=0
4<=x<=9
opens up so it has a low point (6.5, -6.25
parabola
focus | (13/2, -6)=(6.5, -6)
vertex | (13/2, -25/4)=(6.5, -6.25)
semi-axis length | 1/4
focal parameter | 1/2
eccentricity | 1
directrix | y = -13/2